OpenStudy (kittyloli):

A, B, C, and D have the coordinates (-8, 1), (-2, 4), (-3, -1), and (-6, 5), respectively. Which sentence about the points is true? A, B, C, and D lie on the same line. Line AB and Line CD and are perpendicular lines. Line AB and Line CD and are parallel lines. Line AB and Line CD and are intersecting lines, but are not perpendicular. Line AC and Line BD and are parallel lines.

OpenStudy (kittyloli):

@rebeccaxhawaii

OpenStudy (kittyloli):

I don't have a graph, so everything is really difficult

OpenStudy (dustin_whitelock):

I won't give you the answer but here's the hint, the two point-slope formula is y-y1=m(x-x1) where m is the slope of the line, if parallel to each other: m1=m2 if perpendicular, m1m2=-1 Use the eguation to find such relation of m in respect to AB, BC, CD, DA and watch out for the value of m in each case P.S.: the last option- you repeated it by mistake or would it have been something else???

OpenStudy (dustin_whitelock):

BTW, it is not a graphing problem, it should be answered mathematically. I bet this question won't be provided with a graph if you're in the exam hall.

OpenStudy (dustin_whitelock):

@Kittyloli , could you solve it? Or should I do it for you?

OpenStudy (kittyloli):

Fixed

OpenStudy (kittyloli):

I updated the question

OpenStudy (kittyloli):

I'm so confused..

OpenStudy (dustin_whitelock):

@Kittyloli Should I provide you with the solution?

OpenStudy (kittyloli):

Yes please.. :(

OpenStudy (dustin_whitelock):

You have two point slope formula y-y1 = m (x-x1) For line AB, Put x=-8, y=1 (coods. of A) Put x1=-2 ,y1=4 (coods. of B) Put m=m1 Then, 1-4= m1(-8-(-2)) or, -3= m1(-6) or, m1= -3/(-6)= 1/2 For line CD, put x=-3, y=-1; x1=-6, y1=5, m=m2 Now you have, -1-5=m2(-3+6) or, -6=m2(3) or, m2= -2 Here, \[m1\neq m2\] (so, 2nd option wrong) m1m2= -1 So, answer is correct for the 1st choice. P.S.: Now check if the last option is correct or not. :)

OpenStudy (kittyloli):

I think the last one is wrong

OpenStudy (kittyloli):

WAIT, It's right...?

OpenStudy (kittyloli):

@Dustin_Whitelock

OpenStudy (dustin_whitelock):

The last option is wrong too, apparently the first one is correct only. Told you to check the last one too so that you get some practice, :D